### v2

# Yuxia Fu  Zhizhen Zhang11footnotemark: 1  Yuqi Zhang  Zijian Wang  Zi Huang  Yadan Luo

Abstract
Recent Vision-Language-Action (VLA) models reformulate vision-language models by tuning them with millions of robotic demonstrations. While they perform well when fine-tuned for a single embodiment or task family, extending them to multi-skill settings remains challenging: directly merging VLA experts trained on different tasks results in near-zero success rates. This raises a fundamental question: what prevents VLAs from mastering multiple skills within one model? With an empirical decomposition of learnable parameters during VLA fine-tuning, we identify two key sources of non-mergeability: (1) Finetuning drives LoRA adapters in the VLM backbone toward divergent, task-specific directions beyond the capacity of existing merging methods to unify. (2) Action experts develop inter-block dependencies through self-attention feedback, causing task information to spread across layers and preventing modular recombination. To address these challenges, we present MergeVLA, a merging-oriented VLA architecture that preserves mergeability by design. MergeVLA introduces sparsely activated LoRA adapters via task masks to retain consistent parameters and reduce irreconcilable conflicts in the VLM. Its action expert replaces self-attention with cross-attention-only blocks to keep specialization localized and composable. When the task is unknown, it uses a test-time task router to adaptively select the appropriate task mask and expert head from the initial observation, enabling unsupervised task inference. Across LIBERO, LIBERO-Plus, RoboTwin, and multi-task experiments on the real SO101 robotic arm, MergeVLA achieves performance comparable to or even exceeding individually finetuned experts, demonstrating robust generalization across tasks, embodiments, and environments. Project page: https://mergevla.github.io/
1 Introduction
Vision-Language-Action (VLA) models [28, 3, 27, 22, 4, 58, 2, 36, 38, 18, 46, 51, 19] have recently enabled robot agents to perform complex manipulation tasks by fine-tuning large vision-language models (VLMs) with millions of robotic demonstrations. By reformulating action learning as a language generation or policy decoding problem, VLA models inherit broad visual grounding and semantic understanding from VLMs, and have shown notable performance in single-task or single-embodiment settings. However, real-world generalist agents must support multiple skills, embodiments, and environments, requiring the ability to consolidate many independently fine-tuned VLA models into a single unified policy.
A natural approach is model merging, which has proven effective in large language and vision models [21, 48, 31, 8, 53, 20, 39, 54, 56]. Those techniques can integrate multiple specialized models without joint retraining or revisiting their original datasets. Yet, when these approaches are applied to VLA experts fine-tuned on distinct manipulation tasks, the merged model exhibits near-zero success rate. This failure suggests that VLA finetuning induces structural specialization incompatible across tasks, which is rarely observed in conventional VLM merging.
To uncover the root causes, as shown in Figure 1, we perform a fine-grained decomposition of trainable parameters across representative VLA architectures and identify two major sources of non-mergeability: First, LoRA updates within the VLM backbone diverge sharply across tasks. Each manipulation task reshapes the pretrained representation space to satisfy its own perception-control alignment. Direct averaging  [21, 39] or sign-resolved merging [48] thus reactivates irrelevant or even contradictory parameters, corrupting shared vision-language subspaces and degrading task-invariant semantics. Second, the train-from-scratch action decoders accumulate strong task-specific dependencies across blocks through self-attention feedback. This coupling spreads localized task information globally, breaking modularity and preventing compositional merging even under identical architectures and initialization.
Built upon these findings, we introduce MergeVLA, a merge–oriented VLA architecture that preserves mergeability by design. When executing a specific task, MergeVLA applies sparsely activated LoRA adapters, implemented via task masks, to selectively activate the merged parameters contributing to task-relevant responses while suppressing those that mislead other tasks. Moreover, MergeVLA reconfigures the action expert to remove self-attention propagation, relying solely on cross-attention pathways. This eliminates the sustained accumulation of task dependence across blocks, allowing most layers to be effectively merged using simple weight averaging. Due to these architectural modifications, it surprisingly shows strong out-of-distribution (OOD) generalisation by 13.4% higher success rate than VLA-Adapter under varying corruptions. However, the deeper blocks of the action expert, referred to as the expert head, remain unmergeable due to their strong task specialization. Consequently, each task keeps its own expert head, which in practice is typically the final block $L$. To address the challenge of mixed-task evaluation [39, 20], where the task identity is unknown at inference time, we adopt a training-free test-time task router that identifies the most relevant task directly from the input features. For each candidate task, the router runs the VLM with its corresponding task mask to obtain its hidden states, and then measures its response against principal components extracted from the value projections of a shared merged action expert. It selects the task with the highest response and activates the associated task mask and expert head. This allows MergeVLA to generalize to different tasks without additional supervision, enabling a single merged model to adaptively activate the right skill components for execution.
Extensive experiments show that MergeVLA achieves success rates of $\mathbf{90.2\%}$, $\mathbf{62.5\%}$, and $\mathbf{70.7\%}$ on the LIBERO [29], LIBERO-Plus [15], and RoboTwin [7] benchmarks under the mixed-task evaluation setting, and $\mathbf{90.0\%}$ on real-world experiments with the SO101 robotic arm, demonstrating its effectiveness and robustness across cross-skill, cross-environment, and cross-embodiment evaluations. These results show that model merging is not only feasible for VLAs, but can serve as a scalable path toward generalist embodied agents.
2 Related Work
2.1 Vision-Language-Action Models
Recent VLA models leverage the rich commonsense knowledge of large-scale VLMs and are fine-tuned on large-scale robotic trajectory data [43, 35, 4, 58] in an end-to-end manner to obtain more generalizable visuomotor policies [4, 58, 43, 14, 25, 24, 1]. Among them, OpenVLA [28] is a widely used open-source model based on Prismatic-7B [26], which discretizes robot actions into rarely used vocabulary tokens and generates them autoregressively. Similarly, $\pi_{0}$ [3] builds on existing VLMs with a dual-system VLA architecture, introducing a lightweight action expert that uses conditional flow matching to generate continuous action chunks for faster inference. Building on this, $\pi_{0.5}$ [22] retains the text-generation ability of the VLM, enabling simultaneous high-level subtask descriptions and low-level action prediction. Recently, VLA-Adapter [45] proposed a 0.5B-parameter dual-system model that coordinates the VLM and action expert through a novel mechanism, employing chunk-wise autoregressive generation for faster inference and strong performance at a small scale. While existing VLAs perform well on widely used robotic benchmarks [33, 29, 7, 42], their multi-task ability relies on joint training [28, 3, 22, 45], making them inefficient. Besides, their tightly coupled dual systems [55, 9, 5, 3, 45] also hinder model merging.
Figure 1: Comparison between the structures of different VLAs. OpenVLA uses a standard VLM for token-based action generation. VLA-Adapter adds an action expert with cross- and self-attention layers. MergeVLA simplifies this design by removing non-mergeable self-attention layers for effective merging.
Figure 2: Overview of MergeVLA architecture. (1) To address destructive LoRA parameter interference in finetuned VLM, task masks are applied to all merged LoRA modules to selectively activate the merged parameters contributing to task-relevant responses while suppressing those that mislead other tasks. (2) To solve the incompatibility of action experts, the architecture is redesigned to contain only cross-attention blocks and use $\mathrm{sigmoid}$ gate to preserve and rely on robust VLM features. Most blocks then can be merged except deeper blocks named expert head are left unmerged due to their task specification. (3) To address the setting where the task identity is unknown at inference time, a training-free test-time task router is adopted to dynamically select task-specified components by computing task relevance from VLM hidden states in the value-based subspace of the merged action expert.
2.2 Model Merging
Model merging enables efficient knowledge reuse by combining existing model weights to construct a unified model capable of performing multiple tasks. Early studies [16, 34, 47, 23, 32] showed that simple weighted averaging of checkpoints can improve performance and introduce multi-task capability. Task Arithmetic (TA) [21] treats the parameter difference between fine-tuned and pretrained models as a task vector and merges them to integrate task-specific knowledge. Nonetheless, it overlooks various conflicts that may arise among different task vectors. Subsequent methods [48, 53, 10, 12, 52, 40, 8] handle conflicts via parameter pruning or rescaling, whereas subspace-based ones [17, 39, 31, 30] apply low-rank decomposition (e.g., SVD) for more consistent merging. Other studies [56, 54, 39, 57] design merging methods specifically for parameter-efficient fine-tuning (PEFT) modules. In contrast to these pre-merge approaches, methods like [30, 20, 49, 44, 41] perform test-time merging to handle severe task interference, substantially improving performance. Yet, little work has explored model merging in VLA models. The recent ReVLA [11] applies merging to gradually reverse the vision backbone to mitigate visual catastrophic forgetting and enhance domain generalization, rather than to enable multi-task VLA capabilities. To bridge this gap, we propose MergeVLA which enables lightweight multi-task robotic learning.
3 Preliminary
Task Formulation. We consider a collection of single-skill imitation learning datasets $\mathfrak{D}=\{\mathcal{D}_{m}\}^{M}_{m=1}$ where each dataset $\mathcal{D}_{m}$ corresponds to a distinct manipulation task. Each training set is denoted as $\mathcal{D}_{m}=\{\mathbf{I}^{v}_{t},\mathbf{I}^{w}_{t},L\}_{t=1}^{T}$, where $\mathbf{I}^{v}_{t}$ the third-person view image, $\mathbf{I}^{w}_{t}$ the wrist-mounted image, $L$ the task instruction at each time step $t$. Finetuning a pretrained VLA model on each dataset yields $M$ task-specific weights. The objective of model merging is to unify task-specific $\{\Theta_{1},\ldots,\Theta_{M}\}$ into a single general agent $\Theta_{\operatorname{merge}}$ without retraining.
VLA Architectures. Figure 1 illustrates the architectures of different VLAs. Some VLA models [28, 36, 27, 46, 51] directly build upon existing VLMs, where only the language component is non-mergeable as refered to our experiments in Table 1. Others introduce an additional action expert [3, 22, 45], among which VLA-Adapter trains its action expert from scratch. However, this tightly coupled dual-system design makes the overall model difficult to merge. MergeVLA simplifies this structure by removing the non-mergeable self-attention layers, enabling all components except the expert head to be merged effectively.
4 Our Approach: MergeVLA
Finetuning VLA models on individual manipulation tasks, while effective, produces isolated experts that cannot be trivially merged. As discussed in Section 5.2, applying standard model merging approaches to VLA specialists trained on LIBERO [29] results in a complete breakdown, with all merged variants achieving zero success rate. This failure is unexpected given the success of merging strategies in language-only and vision-language domains, and suggests that VLA merging presents a more severe challenge. To understand this phenomenon, we conduct a systematic analysis of the parameter space and architectural behavior of mainstream VLAs. Our findings, as depicted in Figure 3, show that VLA unmergeability arises from two complementary failure modes:
  1. 1.
    Destructive LoRA Parameter Interference: As shown in the left plot, task-specific LoRA updates activate largely disjoint subsets of channels. When merging only four tasks, the proportion of parameters that are relevant to exactly one task (i.e., selfish parameters) already exceeds 75%. Such extreme task exclusivity produces severe parameter conflicts, which directly cause naïve model-merging strategies to fail.
  2. 2.
    Architectural Incompatibility of Action Experts: More critically, resolving LoRA interference is necessary but not sufficient. Even when the VLM is perfectly merged, simply averaging the action experts in architectures such as VLA-Adapter [45] still yields 0% success. The right plot explains why: although the shallow blocks remain moderately aligned across tasks, the parameter distance explodes in the final layers. Such divergence arises because the action expert is trained entirely from scratch and contains self-attention layers that accumulate task-specific differences over depth. Instead of providing modular transformations, these layers propagate and amplify task-dependent signals, causing deeper blocks to become pathologically specialized to individual tasks. As a result, their parameters are inherently irreconcilable under any merging scheme.
Additionally, prior merging studies [39, 20] also note that the mixed-task setting is considerably more challenging than per-task evaluation, where the task identity is unknown. In practice, many approaches rely on hand-picked priors, e.g., task identity or an expert-specific prompt, to perform well, and degrade rapidly without them. Motivated by these insights, we therefore structure MergeVLA as a unified framework that addresses these challenges by: (1) Stabilizing VLM merging via task-specific masking (Sec 4.1) to address extreme LoRA parameter conflict (Q1); (2) redesigning the action expert (Sec 4.2) to mitigate architectural incompatibility (Q2); (3) introducing a learning-free task routing mechanism (Sec 4.3) to enable operation without a task identity at test time (Q3).
4.1 Task Conflicts in LoRA-Finetuned VLM
To address Q1, we first consider a standard LoRA update. Let $\Theta_{0}$ denote the pretrained checkpoint and $\Theta_{m}$ the LoRA-finetuned weights after finetuning on task $m\in\{1,\dots,M\}$. The task vector for task $m$ is defined as $\tau_{m}=\Theta_{m}-\Theta_{0}$. Most data-free methods $\mathcal{R}(\cdot)$ construct a single merged update $\tau_{\mathrm{merge}}$ by merging the updates [21, 48, 20, 17, 39]:
$\tau_{\mathrm{merge}}=\alpha\,\mathcal{R}(\{\tau_{m}\}_{m=1}^{M}),\quad\Theta_{\mathrm{merge}}=\Theta_{0}+\tau_{\mathrm{merge}},$ (1)
where $\alpha$ is the scaling factor. As $\tau_{\mathrm{merge}}$ is unusable directly due to the conflicts we identified, we must move beyond a single global update. We leverage a task-specific binary masking strategy $\mathbf{S}_{m}$ that isolates components in $\Theta_{\mathrm{merge}}$ beneficial to task $m$, while suppressing those encoding conflicts. Formally,
$\Theta_{\mathrm{merge}}^{(m)}=\Theta_{0}+\mathbf{S}_{m}\odot\tau_{\mathrm{merge}},$ (2)
where $\odot$ denotes element-wise multiplication. The mask $\mathbf{S}_{m}$ is constructed by a parameter-level consistency test: A parameter is retained if and only if its task-specific $\tau_{m}$ is both (1) significant and (2) dominant over the scaled residual difference with $\tau_{\mathrm{merge}}$, indicating that it aligns with the overall merge and contributes positively to task $m$:
$\mathbf{S}_{m}=\mathbb{I}\left[\,|\tau_{m}|>\lambda\,|\tau_{\mathrm{merge}}-\tau_{m}|\,\right],$ (3)
where $\mathbb{I}[\cdot]$ is the indicator function that returns $1$ if the condition is true and 0 otherwise, and mask ratio $\lambda$ controls the tolerance for disagreement. This principled approach of pruning parameters based on their alignment consistency, adapted here for LoRA merging, shares its core formulation with task-vector compression methods [44].
Analysis of Parameter Selfishness. Based on this formulation, we compute the proportion of selfish parameters, namely those retained by exactly one task mask:
$\text{ratio}_{\mathrm{selfish}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\!\left[\sum_{m=1}^{M}(\mathbf{S}_{m})_{i}=1\right],$ (4)
Empirically, as Figure 3 (left) shows, this selfish ratio on LIBERO for two merging methods as the number of merged tasks increases from 2 to 4. In all cases, $\text{ratio}_{\mathrm{selfish}}$ rises at around 75%, indicating that most parameters are selfish (kept exclusively by a single task) and underscoring the importance of task-specific masking to mitigate cross-task interference. Notably, applying the task mask encourages some LoRA-finetuned parameters to revert toward their pretrained weights, improving merge stability. A similar effect was observed in ReVLA [11], where VLA finetuning was shown to overwrite pretrained visual knowledge. This forgetting becomes more pronounced when multiple experts are merged, as conflicting task updates distort the shared feature space. By sparsely activating merged LoRA parameters through the proposed mask, MergeVLA preserves pretrained visual–language representations and mitigates cross-task conflicts, enabling more consistent and stable merging across tasks.
Figure 3: Left: Selfish ratio of the masks from TA [21] and TIES [48] by merging different numbers of tasks. The selfish ratio is computed following Equation 4. Right: The average relative L2 distance across blocks between all pairs of action experts.
4.2 Redesigning the Action Expert for Mergeability
Our diagnosis proved that stabilizing the VLM (Q1) is insufficient; the action expert itself is a fundamental barrier (Q2). As shown in Section 5.2, applying only the LoRA task mask still results in 0% success rate. Our analysis pinpoints the incompatibility with the VLA-Adapter [45] architecture (Fig. 1), which consists of $L$ transformer blocks trained from scratch. Each block contains a self-attention layer on the block input $\mathbf{x}^{i}$, two cross-attention layers conditioned on the VLM-provided hidden states $\mathbf{h}_{\text{T}}^{i}$ (task) and $\mathbf{h}_{\text{A}}^{i}$ (action), and a feed-forward network (FFN). A gating function is applied to the task stream, i.e., $\hat{\mathbf{h}}_{\text{T}}^{i}=g(\mathbf{h}_{\text{T}}^{i})$, with $g(\cdot)$ implemented as $\tanh$.
Different from its flawed design, MergeVLA’s redesign (Fig. 2) is a principled response to nonmergeability diagnosis, introducing two key modifications:
  • Remove Self-Attention: We eliminate the self-attention layers, retaining cross-attention only. Since the expert is trained from scratch, self-attention layers develop strong, task-specific biases that are irreconcilable. Removing them forces the expert to rely on the robust, shared VLM features.
  • Replace Gating: We replace the $\tanh$ gate with a $\mathrm{sigmoid}$ gate. The original $\tanh$ gate can suppress VLM signals via negative activations, forcing the expert to rely on its own scratch-trained (and task-specific) parameters. $\mathrm{sigmoid}$ ensures VLM information is always preserved and balanced.
These two architectural changes alone are remarkably effective, achieving an 13.4% higher success rate on the out-of-distribution (OOD) LIBERO-Plus [15] testbed. This confirms our new design better leverages the VLM’s robustness and is inherently more generalizable.
Merging via Specialization Hierarchy. Since the action expert is trained from scratch, existing task vector-based merging approaches are inapplicable here, as there is no shared initialization among different experts. Therefore, we adopt a simple weight-averaging strategy to merge the parameters of the action experts across tasks. We observe that such averaging works surprisingly well for the shallow blocks of the action expert. However, it fails for the deeper blocks, where the parameter discrepancy tasks increase sharply.
This reflects strong task-specific specialization, and we refer to these divergent layers collectively as the expert head, denoted as $\mathbf{H}^{l\rightarrow L}$, spanning blocks $l$ through $L$. In most cases, $l=L$, meaning that only the final block requires separate handling. We hypothesize that under regression-based training objectives, each expert head becomes highly specialized to the action distribution of its corresponding task. Even small discrepancies between these distributions can lead to incompatible weights, making simple parameter averaging ineffective. This effect is particularly pronounced in fine-grained manipulation tasks, where minor output deviations can cause entire trajectories to fail. Therefore, we leave the expert heads unmerged and allow the model to use the corresponding head for each task.
4.3 Test-time Task Routing
When the task identity is known, MergeVLA can already accomplish each skill by manually selecting the corresponding task mask $\mathbf{S}_{m}$ and expert head $\mathbf{H}_{m}^{l\rightarrow L}$. However, to operate without a known task identity (Q3) at inference time, the model must dynamically select these components based solely on the input observations, enabling cross-skill ability at a joint-task level [39]. To this end, we propose a test-time task routing mechanism that infers task relevance directly from the model’s internal parameter subspaces, inspired by recent observations that fine-tuning pushes different tasks into distinguishable parameter subspaces [41]. Given the merged LoRA parameters $\tau_{\mathrm{merge}}$ and task masks $\{\mathbf{S}_{m}\}_{m=1}^{M}$, we obtain $M$ masked VLM variants by applying each mask to the merged weights:
$\Theta^{(m)}_{\mathrm{merge}}=\Theta_{0}+\mathbf{S}_{m}\odot\tau_{\mathrm{merge}},\quad m=1,\dots,M.$ (5)
Each masked VLM produces hidden states $\left[\mathbf{h}_{\mathrm{T}}^{l},\mathbf{h}_{\mathrm{A}}^{l}\right]$ from block $l$, which is then forwarded to the corresponding $l$-th block of the action expert. Inside this block, two cross-attention paths are present: one conditioned on the task hidden state $\mathbf{h}_{\mathrm{T}}^{\,l}$ with parameters $(\mathbf{Q}_{\mathrm{T}}^{\,l},\mathbf{K}_{\mathrm{T}}^{\,l},\mathbf{V}_{\mathrm{T}}^{\,l})$, and the other conditioned on the action hidden state $\mathbf{h}_{\mathrm{A}}^{\,l}$ with $(\mathbf{Q}_{\mathrm{A}}^{\,l},\mathbf{K}_{\mathrm{A}}^{\,l},\mathbf{V}_{\mathrm{A}}^{\,l})$.
A key design choice is which subspace to use for routing. We hypothesize that query $\mathbf{Q}$ and $\mathbf{K}$ govern attentional selection, thus being sensitive to input scaling and risk collapsing into task-specific subspaces. Empirically, value-based subspaces provide more stable and discriminative signals for routing, as they directly encode the task-dependent information written into the hidden states. We therefore analyze the principal components of the value projection matrices V of the $l$-th block for both paths via singular value decomposition (SVD):
$\mathbf{V}_{\mathrm{T}}^{\,l}$ $=\mathbf{L}_{\mathrm{T}}^{\,l}\,\mathbf{\Sigma}_{\mathrm{T}}^{\,l}\,(\mathbf{R}_{\mathrm{T}}^{\,l})^{\top},$ (6)
$\mathbf{V}_{\mathrm{A}}^{\,l}$ $=\mathbf{L}_{\mathrm{A}}^{\,l}\,\mathbf{\Sigma}_{\mathrm{A}}^{\,l}\,(\mathbf{R}_{\mathrm{A}}^{\,l})^{\top}.$
We retain the top-$k_{r}$ right singular vectors in each path to form the dominant content components of the expert subspace: $\mathbf{P}_{\mathrm{T}}^{\,l}\in\mathbb{R}^{k_{r}\times d},$ $\mathbf{P}_{\mathrm{A}}^{\,l}\in\mathbb{R}^{k_{r}\times d},$ formed by taking the first $k_{r}$ rows of $(\mathbf{R}_{\mathrm{T}}^{\,l})^{\top}$ and $(\mathbf{R}_{\mathrm{A}}^{\,l})^{\top}$, respectively. For each task $m$, the hidden state from its masked VLM is projected onto these two subspaces to measure its activation strength,
$r_{\mathrm{T},m}=\big\|\mathbf{P}_{\mathrm{T}}^{\,l}\mathbf{h}^{\,l}_{\mathrm{A},m}\big\|_{2},r_{\mathrm{A},m}=\big\|\mathbf{P}_{\mathrm{A}}^{\,l}\mathbf{h}^{\,l}_{\mathrm{T},m}\big\|_{2}.$ (7)
Let $\mathbf{r}_{\mathrm{T}},\mathbf{r}_{\mathrm{A}}\in\mathbb{R}^{M}$ collect the scores $\{r_{\mathrm{T},m}\}_{m=1}^{M}$ and $\{r_{\mathrm{A},m}\}_{m=1}^{M}$, respectively. Here, the combined score vector is given by $\mathbf{r}=\tfrac{1}{2}\big(\mathbf{r}_{\mathrm{T}}+\mathbf{r}_{\mathrm{A}}\big)$. The routing probabilities are calculated by softmax $p_{m}=\operatorname{softmax}(\mathbf{r})_{m}.$ Then we can use $\arg\max_{m}\,p_{m}$ to choose the task index $m^{*}$. Once $m^{*}$ is determined, the model uses the corresponding task mask $\mathbf{S}_{m^{*}}$ and expert head $\text{H}^{l\rightarrow L}_{m^{*}}$ to perform the forward pass. This design allows the router to infer the underlying task purely from the input-driven hidden representations of the masked VLMs, without any additional training or supervision. In practice, we find that a single routing step using the initial observation at t=0t{=}0 is sufficient to identify the correct task. The selected task mask and expert head are then fixed for the rest of the episode, avoiding repeated routing during inference. Although this mechanism requires maintaining $M$ task masks and their corresponding action heads, the additional computational and parameter overhead is minimal.
5 Experiments
5.1 Experimental Setup
Simulation Benchmarks. We evaluate MergeVLA on three simulation benchmarks:
- LIBERO [29] is a comprehensive benchmark consisting of four distinct task suites: Spatial, Object, Goal, and Long. Each suite contains 10 tasks with 50 demonstrations for every task. The benchmark enables the assessment of the robot’s multi-skill competence, including spatial relationships, object interactions, and task-specific objectives.
- LIBERO-Plus [15] builds upon the original LIBERO benchmark by introducing seven distinct perturbations, as shown in Figure 4. The benchmark comprises 10,030 tasks, providing diverse settings for evaluating model generalization and robustness under shifts.
- RoboTwin 2.0 [7] serves as a cross-embodiment benchmark for dual-arm manipulation. As shown in Figure 5, we select three embodiments and four tasks for a comprehensive assessment of the cross-embodiment cross-skill ability.
Real-World Robot Experiments. For real-world evaluation, we deploy the SO101 robot within the LeRobot framework, as shown in Figure 6. We design three manipulation tasks: cube picking, cube stacking, and cube pushing. Each task includes 50 human-teleoperated demonstrations used for training. The detailed experimental setup is provided in Section 5.5.
Table 1: LIBERO results across task splits. Comparison between finetuned and merged variants of MergeVLA. All numbers are success rates (%). $\mathbf{S}$ indicates that task masks are used during merging. “Params (B)” denotes the total number of model parameters (in billions) required to evaluate on all four tasks, including the LLM backbone and the action expert. Gray-highlighted rows correspond to per-task finetuned checkpoints evaluated on their own tasks, serving as upper-bound references for model merging.
Method Merge Method Merge Part Params (B) Spatial Object Goal Long Avg.
Single-task Finetuned Model
$\mathrm{OpenVLA}$ [28] - - 7 $\times$ 4 84.7 88.4 79.2 53.7 76.5
$\mathrm{VLA}$-$\mathrm{Adapter}$ [45] - - 0.68 $\times$ 4 99.6 99.6 98.2 96.4 98.5
$\mathrm{MergeVLA}$ - - 0.68 $\times$ 4 98.0 98.6 95.0 95.0 96.7
Merged Model
$\mathrm{OpenVLA}$ TA [21] Vision Backbones 7 $\times$ 4 56.6 58.0 55.6 6.6 44.2
$\mathrm{OpenVLA}$ TA [21] All 7 0.0 0.0 0.0 0.0 0.0
$\mathrm{OpenVLA}$ TA [21] + $\mathbf{S}$ All 7 74.2 82.6 68.8 24.0 62.4
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [21] All 0.68 0.0 0.0 0.0 0.0 0.0
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [21] + $\mathbf{S}$ All 0.68 0.0 0.0 0.0 0.0 0.0
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [21] + $\mathbf{S}$ Except $\mathbf{H}^{L\rightarrow L}$ 0.70 50.2 34.6 0.0 7.4 23.1
$\mathrm{MergeVLA}_{\mathrm{EMR}}$ EMR [20] 96.0 63.2 62.0 40.6 65.5
$\mathrm{MergeVLA_{TSV}}$ TSV [17] + $\mathbf{S}$ 99.4 97.8 74.4 54.8 81.6
$\mathrm{MergeVLA_{KnOTS}}$ KnOTS[39] + $\mathbf{S}$ 96.8 98.8 84.8 71.4 88.0
$\mathrm{MergeVLA_{TA}}$ TA [21] + $\mathbf{S}$ Except $\mathbf{H}^{L\rightarrow L}$ 0.70 98.0 98.8 85.4 76.6 89.7
$\mathrm{MergeVLA_{WUDI}}$ WUDI [8] + $\mathbf{S}$ 97.6 98.2 85.6 78.2 89.9
$\mathrm{MergeVLA_{TIES}}$ TIES [48] + $\mathbf{S}$ 94.8 94.6 91.8 79.4 90.2
Implementation Details. Our vision-language backbone is Qwen2.5-0.5B [50]. By default, we set $l=L$, $k_{r}=8$, mask ratio $\lambda=0.6$, merging scaling factor $\alpha=1$. All finetuning are conducted on a single NVIDIA A6000 Ada GPU (48 GB). Other hyperparameters can be found in Appendix.
5.2 Results on LIBERO
We first reproduce results of OpenVLA [28] and VLA-Adapter [45] on four tasks, as highlighted in grey in Table 1. OpenVLA performs notably worse than the others, with an average success rate of 76.5% and only 53.7% on the long-horizon LIBERO-Long task. Our MergeVLA, modified from VLA-Adapter to be merge-friendly, achieves comparable fine-tuning performance, indicating that our structural changes preserve the original capability.
We have tried evaluating single-task finetuned models on unseen tasks (e.g., testing a Spatial expert on the Object suite), all methods achieve 0% success across tasks. This clearly indicates that existing VLA models lack cross-skill generalization. The lower part of Table 1 reports the merging results. Directly applying TA [21] to OpenVLA fails entirely, while merging only the vision and projector components improves to 44.2%. Adding a task-specific mask further raises the average success rate to 62.4%, though performance remains unbalanced. On VLA-Adapter, tight coupling between the VLM and action expert makes merging difficult—even with masking, unless the final action block is excluded. As shown in the blue-highlighted rows, our MergeVLA consistently outperforms other merging baselines. With TIES [48] and WUDI [8], it achieves balanced results across all four tasks (up to 90.2% average success rate), while remaining lightweight and only 6.5% below fine-tuning performance. These results confirm that MergeVLA enables efficient knowledge reuse and strong multi-task performance with a compact VLA model.
Table 2: Robustness of different models under visual and language shifts on LIBERO-Plus. All results are success rates (%) averaged over 4 task suites. Gray-highlighted rows correspond to per-task finetuned checkpoints evaluated on their own tasks, serving as upper-bound references for model merging. Shift definitions: S1 – Background Textures; S2 – Camera Viewpoints; S3 – Language Instructions; S4 – Lighting Conditions; S5 – Object Layout; S6 – Robot States; S7 – Sensor Noise.
Method S1 S2 S3 S4 S5 S6 S7 Avg.
Single-task Finetuned Model
$\mathrm{OpenVLA}$ [28] 34.8 0.8 23.0 8.1 28.5 3.5 15.2 16.3
$\mathrm{\pi_{0}}$ [3] 81.4 13.8 58.8 85.0 68.9 6.9 79.0 56.3
$\mathrm{VLA}$-$\mathrm{Adapter}$ [45] 76.6 36.4 73.8 71.0 70.2 37.4 57.2 59.0
$\mathrm{MergeVLA}$ 92.7 62.4 75.7 92.7 73.7 46.4 74.7 72.4
Merged Model
$\mathrm{VLA}$-$\mathrm{Adapter}$ [45] 15.7 6.6 17.6 11.2 15.0 4.1 7.1 10.8
$\mathrm{MergeVLA}_{\mathrm{TSV}}$ 64.3 44.8 58.8 72.9 59.4 30.4 52.7 53.5
$\mathrm{MergeVLA}_{\mathrm{TA}}$ 78.2 53.1 68.4 79.0 65.0 34.6 62.7 61.6
$\mathrm{MergeVLA}_{\mathrm{TIES}}$ 85.7 50.7 66.0 84.2 68.1 30.3 66.0 62.5
Table 3: RoboTwin success rates (%) of different variants of MergeVLA across embodiments and tasks. $\mathbf{T}_{1}$: Place Container Plate, $\mathbf{T}_{2}$: Handover Block, $\mathbf{T}_{3}$: Open Microwave. Gray-highlighted rows correspond to per-task finetuned checkpoints evaluated on their own tasks, serving as upper-bound references for model merging.
Setting A: Cross embodiments, Single task
Method $\mathbf{T}_{1}$ Avg.
Aloha ARX Piper
Single-task Finetuned 90.0 90.0 84.0 88.0
$\mathrm{MergeVLA}_{\mathrm{TA},\mathbf{H}^{(L-1)\rightarrow L}}$ 86.0 82.0 68.0 78.7
$\mathrm{MergeVLA}_{\mathrm{TIES},\mathbf{H}^{(L-1)\rightarrow L}}$ 88.0 92.0 86.0 88.7
Setting B: Cross embodiments, Cross task
Method $\mathbf{T}_{1}$ $\mathbf{T}_{2}$ $\mathbf{T}_{3}$ Avg.
Aloha ARX Piper
Single-task Finetuned 90.0 46.0 92.0 76.0
$\mathrm{MergeVLA}_{\mathrm{TA},\mathbf{H}^{(L-1)\rightarrow L}}$ 80.0 0.0 66.0 48.7
$\mathrm{MergeVLA}_{\mathrm{TA},\mathbf{H}^{(L-2)\rightarrow L}}$ 82.0 0.0 66.0 49.3
$\mathrm{MergeVLA}_{\mathrm{TIES},\mathbf{H}^{(L-1)\rightarrow L}}$ 90.0 0.0 88.0 59.3
$\mathrm{MergeVLA}_{\mathrm{TIES},\mathbf{H}^{(L-2)\rightarrow L}}$ 88.0 38.0 86.0 70.7
Figure 4: Seven perturbation types in the LIBERO-Plus benchmark, used to evaluate robustness under visual and language shifts.
Figure 5: Experimental setup in the RoboTwin environment, featuring three robotic embodiments and a suite of manipulation tasks for cross-embodiment evaluation.
5.3 Results on LIBERO-Plus
Experiments on the LIBERO benchmark verify that MergeVLA achieves strong cross-task performance within known environments. To further examine the model’s generalization and robustness to unseen scenes, we conduct additional evaluations on LIBERO-Plus, a variant of LIBERO that introduces controlled distribution shifts in both visual appearance and language descriptions. We use the same models trained on LIBERO and directly test them on the shifted LIBERO-Plus environments without any additional finetuning. This setting allows us to rigorously assess MergeVLA’s ability to maintain performance under out-of-distribution visual and linguistic conditions.
From Table 2, we first observe that when using single-task finetuned checkpoints evaluated on their own tasks, MergeVLA exhibits stronger robustness under various perturbations than existing VLA models. This improvement stems from its architecture design, which preserves the pretrained VLM’s inherent robustness to visual and language perturbations. Secondly, in the model-merging setting, we find that applying different merging methods with MergeVLA maintains similar robustness even under cross-task evaluation. Notably, when using TA and TIES merging, the merged model even surpasses OpenVLA, $\mathrm{\pi_{0}}$, and VLA-Adapter in the single-task setting, demonstrating that our MergeVLA can effectively transfer and preserve robustness across tasks.
5.4 Results on RoboTwin
While evaluations on the LIBERO and LIBERO-Plus benchmarks demonstrate that MergeVLA attains high performance and robustness in cross-task settings, these tests are limited to a single embodiment. To examine the effectiveness of MergeVLA under cross-embodiment merging, we selected RoboTwin-2.0 as our evaluation platform, since it supports multiple dual-arm robots and enables a broad assessment of generalization across hardware. Specifically, we designed two experimental settings: A: Three dual-arm robots, Aloha-Agilex, ARX-X5, and Piper, each perform the same manipulation task place container plate. B: The same three robots each perform a different task: place container plate, handover block, and open microwave, respectively. For each combination of {embodiment,task}, we collected 50 demonstration trajectories for finetuning. We report the success rate of (i) single-{task, embodiment} fine-tuned models, and (ii) merged models obtained using TA and TIES strategies under the MergeVLA framework. The detailed results are presented in Table 3, where each result is the success rate over 50 trials. Unlike the experiments on LIBERO and LIBERO-Plus, merging across different embodiments in RoboTwin poses a greater challenge for the test-time task router. We find that only keeping the final block $L$ as expert head is insufficient, as embodiment-specific differences in morphology and action space introduce stronger specialization and conflicts within the action heads. For Setting A, routing $\mathbf{H}^{(L-1)\rightarrow L}$ preserves the performance of individually finetuned policies, indicating that the earlier merged blocks still capture transferable knowledge across different embodiments. For Setting B, routing $\mathbf{H}^{(L-2)\rightarrow L}$ and TIES merging are need to maintain comparable performance, especially for the Handover Block task, which requires coordinated dual-arm motion and thus induces stronger conflicts in the action space. Overall, these findings highlight MergeVLA’s ability for cross-embodiment generalization.
Figure 6: Setup of the real-world SO-101 arm experiments with three cube manipulation tasks.
5.5 Real-World Experiments
Tasks. As shown in Figure 6, we evaluate MergeVLA on three cube-based manipulation tasks using a real SO-101 robotic arm:
(i) Pick & Place: the robot must grasp a cube and place it into a black box; success is recorded when the cube is stably placed inside the container.
(ii) Push Cube: the robot must push the cube into a designated white goal zone; success requires the cube to fully enter the region.
(iii) Stack Cube: the robot must pick up the red cube and place it on top of a blue cube; success is defined by a stable, non-slipping stacked configuration.
Data Collection. We collect demonstrations using the SO-101 arm under a leader–follower teleoperation setup, with two RGB camera views: a fixed top-down camera and a wrist-mounted camera. For each task, we collect 50 demonstrations at a frequency of 20 Hz, with randomized cube starting positions. For the Pick & Place and Push Cube tasks, only the red cube is used during data collection. Each demonstration includes synchronized RGB observations, 6 DoF joint actions, and task instructions. We train MergeVLA for 30k steps per task.
Evaluation Protocol. For each model to be evaluated, we perform 20 rollouts per task, with randomized cube initial positions in every rollout. For Pick & Place and Push Cube tasks, we use cubes with randomly different colors that are unseen in the training data, providing a visual shift evaluation. Success is determined according to the task-specific criteria defined above. We report the success rate as the percentage of successful trials out of the 20 rollouts.
Results. In Table 4, we present both fine-tuning and model-merging results of MergeVLA on the real SO-101 robotic arm. For fine-tuning, MergeVLA achieves high success rates across all three cube-manipulation tasks. Notably, in Pick & Place and Push Cube, the robot is required to operate on cubes whose colors differ from those seen during training. MergeVLA remains robust under this distribution shift, reliably detecting the target object and executing the required manipulation, which highlights its strong visual out-of-distribution generalization in real-world settings.
For model merging, we evaluate MergeVLA using TA and TIES as the merging strategies. TIES-based merging delivers the best overall performance, often matching the results of the corresponding single-task models. This demonstrates that MergeVLA preserves cross-task merging ability even when deployed on physical hardware, and is able to reuse skill components without degradation, an encouraging indication of its practicality for multi-skill real robot systems.
Table 4: Real-world SO-101 robot performance, reported as success rates (%) over 20 rollouts per task.
Method Pick & Place Push Cube Stack Cube Avg.
Single-task finetune 90.0 85.0 95.0 90.0
$\mathrm{MergeVLA}_{\mathrm{TA}}$ 70.0 70.0 60.0 66.7
$\mathrm{MergeVLA}_{\mathrm{TIES}}$ 90.0 90.0 90.0 90.0
5.6 Ablation Study
Impact of Mask Ratio. We analyze the effect of the task mask under different $\lambda$, which controls the active ratio of the mask. We vary $\lambda$ from 0.2 to 0.9 and obtain the merged task vector using Task Arithmetic (TA). Figure 7 (a) shows the active ratio of the mask across four tasks. As $\lambda$ increases, fewer parameters remain active, with the LIBERO-Spatial task consistently showing the highest activation ratio, indicating its dominant weight contribution. We further evaluate the impact of $\lambda$ on the model’s performance in the LIBERO-Long task, as shown in Figure 7 (b). When $\lambda$ is small (e.g., 0.2), the mask activates too many parameters, leading to severe task interference and even complete failure. In contrast, when $\lambda$ lies between 0.6 and 0.9, the success rate exceeds 70%. These results suggest that moderate sparsity enables an effective balance between task-specific and merged task vectors. When the parameter differences across tasks are large, relying more on the pretrained model yields better results, whereas emphasizing the merged task vector becomes effective only when the task-specific weights dominate.
Impact of the Subspace Used for Routing. To validate our task router design, we experiment on all four LIBERO task suites using three configurations: (1) using $\mathbf{K}$ projections, (2) using $\mathbf{V}$ projections, and (3) using $\mathbf{K}$ and $\mathbf{V}$ projections jointly, while fixing $\lambda=0.6$ for the task-specific mask and adopting TA for VLM merging. The results are summarized in Table 5. We observe that for Spatial and Long tasks, the choice of projection combination has little effect on performance. Yet, for Object and Goal, using only $\mathbf{K}$ or combining $\mathbf{K}$ and $\mathbf{V}$ leads to a dramatic drop in success rate, and even complete failure in some cases. By inspecting the router’s task selection, we find that when using $\mathbf{K}$, the router tends to misassign tasks. From the perspective of attention interaction, the value projection captures the actual behavioral semantics retrieved by the query, whereas the key projection primarily defines the similarity structure of the query. Consequently, value-based alignment provides a more reliable indicator of task identity.
Figure 7: The ablation study and analysis of $\lambda$. (a) The mask active ratio across different $\lambda$ of LIBERO. (b) The success rate across different $\lambda$ of LIBERO-Long.
Table 5: Ablation results of MergeVLA with different subspaces used for routing on LIBERO.
Method Spatial Object Goal Long Avg.
Only $\mathbf{K}$ 98.0 0.0 39.6 76.6 53.6
$\mathbf{K}\&\mathbf{V}$ 98.0 0.0 85.8 76.6 65.1
Only $\mathbf{V}$ 98.0 98.8 85.4 76.6 89.7
6 Conclusion
The lack of cross-skill capability in existing VLA models remains a critical yet unexplored challenge. We show that current model merging techniques cannot be directly applied to VLAs due to destructive LoRA parameter interference and architectural incompatibility of action experts. We propose MergeVLA, a framework for merging VLA models that is compatible with mainstream merging methods toward an embodied generalist. Experiments across three benchmarks demonstrate MergeVLA’s strong multi-task ability and robustness under distribution shifts, with transferability across embodiments. Future work will explore this promising direction further, including whether larger VLM backbones remain compatible with our framework and whether pretraining on diverse robot datasets can further enhance merging effectiveness.
7 Acknowledgments
This work was partially supported by ARC DE240100105, DP240101814, DP230101196, BA24006, and ARC Industrial Transformation Research Hubs IH230100013.
Supplementary Material
8 Experimental Details
We summarize all fine-tuning hyperparameters used across LIBERO, RoboTwin, and real-world SO-101 experiments in Table 6. All experiments use the same VLM backbone and training configuration unless otherwise specified.
Table 6: Fine-tuning hyperparameters used in all experiments.
Hyperparameter Value
Backbone Qwen-2.5 (0.5B)
Batch size 8
Learning rate $5\times 10^{-4}$
LoRA rank 32
Use proprioception True
Num images 2 (3 for Robotwin)
Gradient step 30k (50k for LIBERO-Long)
9 Algorithm Details
In this section, we give a detailed algorithm description in Algorithm. 1 of how MergeVLA performs inference using our test-time task router when the task identity is unknown.
Algorithm 1 Test-time Task Routing and Inference in MergeVLA
1:Inputs: Task masks $\{\mathbf{S}_{m}\}_{m=1}^{M}$; Expert heads $\{\mathbf{H}_{m}^{\,l\rightarrow L}\}_{m=1}^{M}$; Pretrained VLM weights $\Theta_{0}$; Merged task vector $\tau_{\mathrm{merge}}$; Value projections of the action expert at block $l$: $\mathbf{V}_{\mathrm{T},m}^{\,l},\;\mathbf{V}_{\mathrm{A},m}^{\,l}$; Initial observation $(\mathbf{I}_{0}^{v},\mathbf{I}_{0}^{w},L)$
2:Routing phase (at $t=0$):
3:for $m=1$ to $M$ do
4:   $\Theta_{\mathrm{VLM}}^{(m)}=\Theta_{0}+\mathbf{S}_{m}\odot\tau_{\mathrm{merge}}$ $\triangleright$ Construct masked VLM
5:   $\mathbf{h}_{\mathrm{T},m}^{\,l},\;\mathbf{h}_{\mathrm{A},m}^{\,l}=\Theta_{\mathrm{VLM}}^{(m)}(\mathbf{I}_{0}^{v},\mathbf{I}_{0}^{w},L)$ $\triangleright$ Extract $l$-th block hidden states
6:   $\mathbf{V}_{\mathrm{T},m}^{\,l}=\mathbf{L}_{\mathrm{T},m}^{\,l}\,\mathbf{\Sigma}_{\mathrm{T},m}^{\,l}\,(\mathbf{R}_{\mathrm{T},m}^{\,l})^{\top}$
7:   $\mathbf{V}_{\mathrm{A},m}^{\,l}=\mathbf{L}_{\mathrm{A},m}^{\,l}\,\mathbf{\Sigma}_{\mathrm{A},m}^{\,l}\,(\mathbf{R}_{\mathrm{A},m}^{\,l})^{\top}$
8:   $\mathbf{r}_{\mathrm{T},m}=\big\|\mathbf{P}_{\mathrm{T},m}^{\,l}\mathbf{h}^{\,l}_{\mathrm{A},m}\big\|_{2}$ $\triangleright$ Choose top-$r_{k}$ singular vectors from $\mathbf{R}$ to get $\mathbf{P}$
9:   $\mathbf{r}_{\mathrm{A},m}=\big\|\mathbf{P}_{\mathrm{A},m}^{\,l}\mathbf{h}^{\,l}_{\mathrm{T},m}\big\|_{2}$
10:   $\mathbf{r}_{m}=\tfrac{1}{2}\big(\mathbf{r}_{\mathrm{T},m}+\mathbf{r}_{\mathrm{A},m}\big)$
11:end for
[…]
13:return $m^{*}$
14:Inference phase: Use $\mathbf{S}_{m^{*}}$, and expert head $\mathbf{H}_{m^{*}}^{\,l\rightarrow L}$ for all $t\geq 0$.
10 Preliminary Investigation on OpenVLA
In the early stage of this work, we explored the feasibility of directly applying existing model merging methods to OpenVLA [28], a popular VLA model. OpenVLA consists of three main components: a vision backbone, a projector, and a language model. The language model itself contains 32 transformer blocks followed by a single-layer MLP head (lm_head). We first attempted to merge all components of OpenVLA using Weighted Average and Task Arithmetic methods. However, the merged checkpoint completely failed on all tasks. This was surprising, as OpenVLA is essentially a VLM, and previous studies [13, 6, 37] have shown that VLMs can usually be merged successfully. This prompted us to investigate which part of OpenVLA prevents successful merging.
10.1 Non-mergeable Components in OpenVLA
To locate the source of failure, we decomposed the model into four submodules: $\mathrm{A.}$ the vision backbone; $\mathrm{B.}$ the projector; $\mathrm{C.}$ the language model body (excluding lm_head); and $\mathrm{D.}$ the lm_head. We then merged each submodule separately across the four official LIBERO task checkpoints using the existing merging method Iso-CTS [31]. Each merged checkpoint was evaluated on 50 trials per subtask. The results are summarized in Table 7, where the gray-highlighted row denotes the single-task fine-tuning performance. From the table, we observe that merging modules $\mathrm{A}$, $\mathrm{B}$, or $\mathrm{D}$ only slightly decreases success rates, whereas merging the language model body ($\mathrm{C}$) causes complete failure on all tasks. This clearly indicates that the language model is the primary source of merging failure.
We hypothesize that this phenomenon arises because VLA tasks impose much stricter precision requirements on the model outputs than typical LLM or VLM tasks. In LLMs or VLMs, outputs are often discrete token sequences or probability distributions, where small deviations are tolerable. In contrast, robotic control requires continuous numeric outputs, where even minor errors can cause irreversible physical or simulated state changes. Once the environment diverges from the model’s training distribution, subsequent actions fail catastrophically. As component $\mathrm{C}$ is directly responsible for decoding actions, it likely accumulates task-specific differences that make naive merging infeasible. This also explains why, as shown in the main paper, applying task masks to preserve localized task information can effectively mitigate such conflicts and enable multi-task unification.
Additional patterns can be observed from Table 7. Interestingly, merging only the vision backbone ($\mathrm{A}$) consistently yields higher success rates than merging both A and B together. This counterintuitive result suggests that, in robotic domains, all modules may exhibit nontrivial task interference, and increasing the number of merged modules amplifies this conflict. In contrast, merging the lm_head ($\mathrm{D}$) has little impact on performance. It is only about 3 points below the fine-tuned baseline. Moreover, combinations such as $\mathrm{A+D}$ and $\mathrm{A+B+D}$ show negligible difference from $\mathrm{A}$ and $\mathrm{A+B}$ respectively. To further confirm this, we swapped the lm_head ($\mathrm{D}$) between the LIBERO-Object and LIBERO-Spatial tasks and tested the model on LIBERO-Spatial. Remarkably, it still achieved an 82% success rate over 20 trials, indicating that heads of OpenVLA are largely interchangeable across tasks.
Table 7: Success rates on the four LIBERO task suites when merging different components of OpenVLA [28] using the Iso-CTS [31] merging method. Each merged checkpoint combines four task-specific models, while unmerged components retain their original weights. During evaluation, each subtask is tested with 50 trials. Gray-highlighted row indicates the success rates of individually fine-tuned models.
Method Spatial Object Goal Long Avg.
Finetuned 84.7 88.4 79.2 53.7 76.5
$\mathrm{A}$ 61.4 60.8 59.0 8.4 47.4
$\mathrm{A+B}$ 56.6 58.0 55.6 6.6 44.2
$\mathrm{C}$ 0.0 0.0 0.0 0.0 0.0
$\mathrm{D}$ 83.4 88.8 72.6 49.6 73.6
$\mathrm{A+D}$ 61.0 61.0 62.6 8.4 48.3
$\mathrm{A+B+D}$ 58.0 57.4 53.8 7.4 44.2
10.2 Progressive Block-wise Merging of the Language Model
Figure 8: Success rate on the LIBERO-Spatial task when progressively merging the first $k$ language model blocks of OpenVLA [28] using the Iso-CTS [31] merging algorithm. Each configuration merges four task-specific checkpoints and is evaluated over 10 trials per subtask.
To further analyze why the language model component cannot be merged, we conducted a block-wise study by progressively merging the first $k$ transformer blocks (from 1 to 32) while keeping all other parts fixed to the LIBERO-Spatial task weights. Each merged model was evaluated on 10 trials per subtask, and results are shown in Figure 8. When merging only a few shallow blocks (e.g., up to 8), the model still achieved roughly 80% success rate. However, as the number of merged blocks increased, performance degraded sharply, and beyond 21 merged blocks the model completely failed. This again validates our hypothesis: Task conflicts grow with layer depth, and deeper layers show stronger task-specific divergence that hinders effective merging.
Figure 9: Mask active ratios for each LIBERO task suite, computed for both the vision backbone and the language model components following the same definition as in the main paper. Masks are obtained using the Task Arithmetic [21] merging method with $\lambda=0.6$.
11 Visualization of Mask Ratios in the Vision Backbone and Language Model
To examine how task masks behave across different components of the VLM, we visualize the mask active ratio for each LIBERO task suite. The mask active ratio measures the proportion of positions where the task mask is active (i.e., set to True), indicating that the model uses the pretrained weight + task vector at that location. In contrast, inactive positions fall back to the pretrained weights only. Because the VLM consists of a vision backbone and a language model, we compute the active ratio separately for these two parts to analyze their task-specific behavior. A higher active ratio suggests stronger task-specific contributions, while a lower ratio indicates greater reliance on pretrained weights. Figure 9 shows that the patterns across tasks differ markedly between the vision backbone and the language model. For example, LIBERO-Long exhibits very low activation in the vision backbone but the highest activation in the language model, whereas LIBERO-Object shows the opposite trend—high activation in the vision backbone but minimal activation in the language model. LIBERO-Spatial, in contrast, maintains relatively high and balanced activation across both components. These observations suggest that visual and linguistic pathways contribute task-specific information in distinct ways, offering useful insights for future work on understanding and leveraging task specialization in VLA models.